OpenStudy (anonymous):

What is the theoretical probability of landing on heads if you flip a coin 100 times and land on heads 28 times?

OpenStudy (anonymous):

Do the same for 31, and 41.

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

I'm not that good at probability

OpenStudy (anonymous):

Ok but can you try? Nobody has helped me and I've been at this for almost 2 hours :(

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@~Samson245~

OpenStudy (anonymous):

@DarkMoonZ @camerondoherty

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

sowwy i suck at probability :(

OpenStudy (anonymous):

Ok :']

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

I read your question a couple of times and I can't make of sense of it. What is the probability of landing on heads 100 times and landing on heads 28 times?

OpenStudy (anonymous):

no, what is the theoretical probability of landing on heads 28 times when you flip a coin 100 times

OpenStudy (anonymous):

oh yeah i messed up in d question

OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!! I DONT HAVE TIME AND THIS IS DRIVING ME CRAZY!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

@CallMeKiki

OpenStudy (freckles):

Ok that makes a little more sense say we wanted to know the probability of landing heads 1 times when we flip the coin say 3 times... Well if we flip the coin 3 times these are the possibilities: TTT TTH THH THT HTH HTT HHH HHT There are 8 possibilitites but only 3 of these have H 1 times. So the probability of getting 1 heads given 3 tosses would be 3/8 So the probability of getting 2 heads given 3 tosses would be 3/8 So the probability of getting 0 heads given 3 tosses would be 1/8 The probability of getting 3 heads gives 3 tosses would be 1/8 So hmm... how can we get the denominator for your problem... Well for the example the 8 comes from doing 2^3 The 3 comes from the number of tosses and the 2 comes from the possible head or tails outcome. So I think you can find your denominator pretty easily now. As for the numerator... Well let's think about the previous example what is 3 choose 1 what is 3 choose 2 what is 3 choose 0 what is 3 choose 3

OpenStudy (anonymous):

choose?

OpenStudy (freckles):

\[n \text{ choose k} \text{ is the same as saying } \left(\begin{matrix}n \\ k\end{matrix}\right)\]

OpenStudy (freckles):

So you could develop a formula for this Suppose 0<k<n where k and n are integers of course. The probability of getting k heads given n tosses is: \[\frac{\left(\begin{matrix}n \\ k\end{matrix}\right)}{2^n}\]

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

how do u solve?

OpenStudy (freckles):

I gave you formula try to use it

OpenStudy (freckles):

what is your k and what is your n given

OpenStudy (anonymous):

hmm.....

OpenStudy (anonymous):

i dont know for real.... im REALLY not good at math :(

OpenStudy (freckles):

"The probability of getting k heads given n tosses is:" Going back to part of the formula I gave... So you don't know how to determine what your k and n is from this? you don't know how many tosses you have or the number of heads you want?

OpenStudy (anonymous):

oh 28/100

OpenStudy (freckles):

no no... I'm asking what is k and what is n k i 28 and n is 100 so go back to the rest of the formula I gave and enter those in

OpenStudy (anonymous):

yup, n 28 k 100?

OpenStudy (freckles):

The probability of getting k heads given n tosses is: \[\frac{\left(\begin{matrix}n \\ k\end{matrix}\right)}{2^n} \] you have: The probability of getting 28 heads given 100 tosses is \[\frac{\left(\begin{matrix}100 \\ 28\end{matrix}\right)}{2^{100}} \]

OpenStudy (freckles):

I suggest leaving it just like that because both top and bottom are really huge

OpenStudy (anonymous):

haha ok so then can you tell me the final answer?

OpenStudy (freckles):

that is the final answer

OpenStudy (anonymous):

ok is it (100/28) / 2 (tiny)100 ?

OpenStudy (anonymous):

is that the answer?

OpenStudy (freckles):

no

OpenStudy (freckles):

100 choose 28 is not the same as 100/28

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

can you also help with these? http://openstudy.com/users/77777jeannie77777#/updates/54f76bdee4b0c8e441be8634

OpenStudy (anonymous):

look at the document

OpenStudy (anonymous):

please i need help with #3, #5

OpenStudy (freckles):

you know you are flipping two coins not 1

OpenStudy (anonymous):

yeah i said that right?

OpenStudy (freckles):

for this problem we only flipped one coin

OpenStudy (anonymous):

AHHH OH NO !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (freckles):

flipped one coin hundred times that is

OpenStudy (anonymous):

omg can u pls just tell me the answer

OpenStudy (freckles):

anyways i might help later i must go

OpenStudy (anonymous):

no i already flipped 2

OpenStudy (anonymous):

please!!! i barely have time :'(

OpenStudy (anonymous):

i just forgot to mention that it was 2

OpenStudy (anonymous):

@inowalst

OpenStudy (freckles):

@77777jeannie77777 just so you know the question you asked here on this page is way more advanced than any question given to you on your paper (that is probably why you didn't know what n choose k meant) I think if you follow what @directrix said you have what you really seek you asked me what is the Probability of getting 28 heads given 100 tosses... None of your questions asked that What is probability of getting two heads given a coin (where you have two coins) toss was already answered actually... Follow what @directrix has given you.